Molybdenum has an atomic raduis of 0.145nm. The volume of its cubic unit cell is 0.0375 nm^3 . what is the geometry of the molybdenum unit cell?
1
Expert's answer
2020-11-17T10:40:27-0500

that. First, a=V1/3

V=0.0375; therefore,

a=(0.03751/3, then

0.03751/3*1.732/4= 0.145 and that is the radius given in the problem. sqrt 3 = 1.732